# Learning or Self-aligning? Rethinking Instruction Fine-tuning

### v2

@@ Abstract
Instruction Fine-tuning (IFT) is a critical phase in building large language models (LLMs). Previous works mainly focus on the IFT’s role in the transfer of behavioral norms and the learning of additional world knowledge. However, the understanding of the underlying mechanisms of IFT remains significantly limited. In this paper, we design a knowledge intervention framework to decouple the potential underlying factors of IFT, thereby enabling individual analysis of different factors. Surprisingly, our experiments reveal that attempting to learn additional world knowledge through IFT often struggles to yield positive impacts and can even lead to markedly negative effects. Further, we discover that maintaining internal knowledge consistency before and after IFT is a critical factor for achieving successful IFT. Our findings reveal the underlying mechanisms of IFT and provide robust support for some very recent and potential future works.
@@ 1 Introduction
The advent and evolution of large language models (LLMs) have marked a significant milestone in natural language processing (NLP) (Brown et al., 2020; Touvron et al., 2023a; Touvron et al., 2023b; Zhao et al., 2023a). As one of the core steps in the construction of LLMs, instruction fine-tuning (IFT) employs supervised instruction-response pairs to fine-tune LLMs, thereby facilitating the transformation of LLMs from a continuous writing model to a question-answering agent  (Chung et al., 2022; Iyer et al., 2023; Jang et al., 2023).
Despite the critical role of IFT in the construction of LLMs, there is a significant lack of in-depth research on the mechanisms by which IFT operates. In traditional machine learning, supervised learning aims to fit models to specific tasks and data distributions (Goodfellow et al., 2016;Bishop, 2006), whereas the impact of IFT on LLMs is markedly different. As shown in Figure 1, on one hand, one of the most apparent effects of IFT is its ability to align the output of LLMs more closely with the latent behavioral norms contained within the IFT data, thereby enabling more effective parameter knowledge expression (Zhou et al., 2023;Chen et al., 2023a). On the other hand, many existing studies aim to facilitate domain-specific adaptation of LLMs through IFT, by injecting the world knowledge contained in IFT data into LLMs (Li et al., 2023; Cui et al., 2023;Chen et al., 2023b).
Figure 1: Two potential mechanisms for instruction fine-tuning. 1) Learning, which injects world knowledge in IFT data into the LLMs; 2) Self-aligning, which aligns queries with knowledge already in LLMs with similar behavioral norms. Elements with the same color are related.
@@
To this end, this paper designs a knowledge intervention framework for analyzing the underlying mechanisms of IFT. The main idea of our framework is to control the consistency between the knowledge in IFT data and the existing parameter knowledge of LLM, in order to decouple the injection of domain knowledge from the transfer of behavioral norms during IFT. This allows for a separate analysis of the roles of these two crucial IFT factors. Specifically, we first employ in-context learning (ICL) (Dong et al., 2023;Brown et al., 2020) to probe the internal parameter knowledge of LLMs. Building on this, we intervene in the composition of existing parameter knowledge and the newly introduced world knowledge within IFT data and then observe the differences in model behavior after IFT using different intervention groups. Based on the framework, we conduct an in-depth analysis to answer the following two critical research questions (RQ):
For RQ1, we initially discover that significant discrepancies between the world knowledge contained in IFT data and the existing parameter knowledge within LLM can substantially undermine the model’s capabilities. Performance derived from a set of IFT data that contains incorrect world knowledge but aligns with the model’s parameter knowledge is significantly better than that from a set containing correct world knowledge but inconsistent with the model’s internal parameter knowledge. To dive into this phenomenon, we explicitly supply LLMs with the world knowledge necessary for answering the instruction, integrated into the context. This strategy allows the model to focus on transforming the output behavioral norms instead of jointly learning the inconsistent world knowledge. We discover that the detrimental effects caused by the inconsistency between parameter knowledge and world knowledge can be significantly mitigated by explicitly providing such self-contained IFT data points. These two findings indicate that attempting to introduce world knowledge through IFT that is inconsistent with the model’s parameter knowledge can severely undermine the model, which suggests that the injection of world knowledge does not lie in the center of a successful IFT.
@@
Our experiments reveal the fundamental role of IFT in the construction of LLMs. Essentially, IFT is not a supervised, domain-specific learning process, but a process of self-aligning instruction with existing internal knowledge (e.g., that can be obtained through few-shot in-context learning probing) of LLMs. Our findings not only provide robust theoretical support for the very recently emerged research on self-play Chen et al. (2024) and super-alignment Burns et al. (2023), etc., but also shed light on the future direction of data construction, model learning, and model evaluation for IFT.
@@ 2 Related Work
By observing token distribution shift of models before and after IFT, Lin et al. (2023) found that most shifts occur with stylistic tokens, strongly supporting the superficial alignment hypothesis (Zhou et al., 2023) and related works on IFT data construction (Chen et al., 2023a;Shen, 2024) and proxy-guided decoding (Liu et al., 2024), etc (Gudibande et al., 2023). While providing intuitive insights, they fall short of providing a comprehensive analysis of IFT’s underlying mechanisms.
Meanwhile, recent efforts have focused on achieving alignment autonomously, such as self-align (Sun et al., 2023;Lu et al., 2024;Guo et al., 2024), self-rewarding (Yuan et al., 2024), self-play (Chen et al., 2024) and supervising stronger models with weaker ones Burns et al. (2023). Despite repeated validations of their effectiveness, there is limited understanding of their success.
@@ 3 Knowledge Intervention Framework
Specifically, we select four multi-choice datasets from different domains. For each question in the four datasets, we employ in-context learning to probe the existing internal parameter knowledge of LLMs. Then, we construct multiple datasets of IFT by adjusting the consistency degree of the knowledge within IFT data and model parameter knowledge. Finally, we analyze the underlying impact of different degrees of world knowledge injection on IFT by fine-tuning LLMs under different settings and examining their performance on the homogeneous, in-domain, and out-of-domain test sets. In the following, we will introduce our knowledge intervention framework in detail.
@@ 3.1 Domain IFT Corpus Setup
In order to facilitate more efficient knowledge consistency identification and IFT evaluation, we select all IFT corpus in the form of multiple choice from four domains: medicine, history, engineering, and jurisprudence. For medicine, we craft a dataset with 20,000 training, 2,206 testing, and 10 development instances by filtering MedMCQA (Pal et al., 2022) for entries with explanations and one correct answer, then applying random sampling. For the other three domains, we procure the relevant items from Xiezhi Benchmark (Gu et al., 2023) and held 10 for development and 250 for test  [^1] The availability of multiple-choice datasets tailored to specific domains is extremely limited. Consequently, from this benchmark, we select the three domains with the largest volume of data..
To make a comprehensive evaluation, for each domain’s IFT, we construct three types of test sets: 1) homogeneous test set (HOMO), which is held out from the same multiple-choice dataset of the domain; 2) in-domain test set (ID), including data from MMLU Hendrycks et al. (2021) that belong to the domain; 3) an out-of-domain test set (OOD), comprising instances in MMLU that are from distinct domains. By observing the performance differences across the three types of test sets, we aim to examine the impact of IFT on various scenarios.
@@ 3.2 Parameter Knowledge Probing via Few-shot In-context Learning
Our knowledge intervention framework relies on effectively detecting the parameter knowledge of pre-trained LLMs. To this end, this paper leverages few-shot in-context learning (Dong et al., 2023;Brown et al., 2020), which is a widely-used approach for probing the capabilities and internal knowledge of pre-trained LLMs (Zhang et al., 2023;Wan et al., 2024), to effectively identify the parameter knowledge of our base LLMs. Specifically, we utilize in-context learning to probe the base model’s response to each data item in domain multi-choice dataset and regard the response as the model’s parameter knowledge for this question.
@@ 3.3 Construction of Instruction Data
Upon probing the intrinsic knowledge of pre-trained LLMs, we build instances based on the consistency between world knowledge contained within the domain data and parameter knowledge, thereby constructing different IFT datasets. Specifically, for each domain and base model, we construct three IFT data groups, including:
  • Harmonious setting, which consists of data where the embedded world knowledge is consistent with model parameter knowledge, which means the pre-trained LLM can answer correctly under in-context learning. In the learning process under this setting, the model only needs to transfer behavioral norms, without the need to acquire additional world knowledge due to the above consistency.
@@
  • Self-aligning setting, which consists of data in which the queries are exactly the same as those in the incompatible set, but we modify the answers corresponding to each query to match the pre-trained LLM’s internal knowledge. Therefore, under this setting, all responses are incorrect, and the model cannot learn any additional world knowledge.
To ensure a fair comparison, we maintain the same size across the three groups of data. Meanwhile, to prevent the potential collapse during model training due to the exclusive use of multiple-choice questions, we generate an explanation for the answer and incorporate an equal proportion of general instruction data sampled from alpaca-gpt4-en (Peng et al., 2023), thereby ensuring a more stable IFT.
@@ 3.4 Experiment Setup
We use LLaMA-2-7B, LLaMA-2-13B, LLaMA-2-70B (Touvron et al., 2023b), and Mistral-7B (Jiang et al., 2023) as the base models of our experiments.
@@
We only calculate loss on outputs, setting epoch to 3, learning rate to $2e^{-5}$ and batch size to 256. For Mistral-7B, we set learning rate to $1e^{-5}$  [^2] Training loss of Mistral-7B using learning rate $2e^{-5}$ does not converge and even spikes.. We use DeepSpeed ZeRO3 Rasley et al. (2020) for LLaMA-2-70B and FSDP Zhao et al. (2023b) for the other three. All experiments are implemented on Nvidia A100-80GB GPUs.
@@ 4 Exp-I: Does Learning Domain-specific World Knowledge Matter for IFT?
Table 1: The performance of the four base LLMs after fine-tuning under harmonious (HAR), incompatible (INC), and self-aligning (SELF) settings. For each domain and base model, models fine-tuned on the harmonious dataset and on the self-aligning dataset achieve superior performance compared to those fine-tuned on the incompatible dataset, across all scenarios including homogeneous (HOMO), in-domain (ID), and out-of-domain (OOD) evaluations. The best/second-best performance for each domain and base model in each evaluation is highlighted in bold/underline.
@@
Table 1 reveals that models fine-tuned under the harmonious setting outperform those fine-tuned under the incompatible setting across homogeneous, in-domain, and out-of-domain evaluations. Specifically, the harmonious setting yields mean performance gains of 11.27%, 14.58%, and 14.57% over the incompatible setting for homogeneous, in-domain, and out-of-domain tests, respectively. The results indicate that utilizing IFT data, which is consistent with model parameter knowledge and does not inject any additional domain-specific world knowledge, yields superior fine-tuned models. This conclusion holds true across homogeneous, in-domain, and out-of-domain evaluations.
@@
To further investigate the impact of learning domain-specific world knowledge on IFT, we conduct a more direct comparative experiment between the self-aligning and incompatible settings. For each domain and pre-trained LLM, the two settings’ datasets use identical queries that the model cannot answer correctly under in-context learning but have different responses: the incompatible dataset’s responses are correct, reflecting world knowledge, while the self-aligning dataset’s responses represent model parameter knowledge, are incorrect.
@@
The above results demonstrate a significant decline of performance in models fine-tuned using data that contain correct world knowledge but conflict with model parameter knowledge, compared to using consistent IFT data aligned with parameter knowledge. This suggests that injecting additional world knowledge through IFT, in cases where it is inconsistent with the parameter knowledge, is unlikely to yield the benefits we anticipate. Therefore, the core role of IFT may lie in facilitating the transfer of behavioral norms, rather than injecting additional domain-specific world knowledge. To further validate this conclusion, in the next section, we design a novel method to decouple the conflict knowledge contained in IFT data and present further analysis of this issue.
@@ 5 Exp-II: IFT with Contextualized Knowledge
To this end, we start with the data of the incompatible setting to construct a dataset with contextualized knowledge. Specifically, given an instruction-answer pair in the incompatible group, we employ GPT-3.5333https://openai.com/ to generate the world knowledge that is required to answer this question. The knowledge is then concatenated with the original instruction, as well as the answer, to construct an augmented instance. Finally, we use the constructed data to fine-tune LLMs and compare them with the models fine-tuned with vanilla IFT.
@@
Table 2 shows the results of contextualized knowledge decoupling in three base models. From the table, it can be observed that fine-tuning the model with data using explicit contextualized knowledge significantly mitigates the adverse effects caused by inconsistencies between parameter knowledge and world knowledge in IFT data. Compared to vanilla IFT using incompatible data, our method achieves an average improvement of 8.16% on LLaMA-2-7B, 9.48% on LLaMA-2-13B, and 3.98% on Mistral-7B. Except in rare cases, IFT using data with contextualized knowledge can significantly improve the effect of IFT across homogeneous, in-domain, and out-of-domain evaluations.
The results indicate that we should not force the model to learn additional inconsistent knowledge when the world knowledge in IFT data is inconsistent with model parameter knowledge. Instead, by decoupling knowledge learning and the transfer of behavioral norms, the problems caused by the above knowledge conflicts can be effectively alleviated. This further verifies our observation in Section 4 that additional world knowledge injection may be ineffective or even harmful for IFT. During IFT, the model should focus on transferring the behavioral norm relying on the existing parameter knowledge and the regularity in the IFT data, rather than learning additional world knowledge. Therefore, for RQ1, we conclude:

### v3

@@ Abstract
Instruction Fine-tuning (IFT) is a crucial phase in building large language models (LLMs). Previous works mainly focus on the IFT’s role in the transfer of behavioral norms and the learning of additional world knowledge. However, the understanding of the underlying mechanisms of IFT remains significantly limited. In this paper, we design a knowledge intervention framework to decouple the potential underlying factors of IFT, thereby enabling individual analysis of different factors. Surprisingly, our experiments reveal that attempting to learn additional world knowledge through IFT often struggles to yield positive impacts and can even lead to markedly negative effects. Further, we discover that maintaining internal knowledge consistency before and after IFT is a critical factor for achieving successful IFT. Our findings reveal the underlying mechanisms of IFT and provide robust support for some very recent and potential future works. We release our experimental dataset and codes to facilitate future work  [^1] https://github.com/renmengjie7/Self-Aligning.
@@ 1 Introduction
The advent and evolution of large language models (LLMs) have marked a significant milestone in natural language processing (NLP) Brown et al. (2020); Touvron et al. (2023b, a). As one of the core steps in the construction of LLMs, instruction fine-tuning (IFT) employs supervised instruction-response pairs to fine-tune LLMs, thereby facilitating the transformation of LLM from a continuous writing model to a question-answering agent Chung et al. (2024); Iyer et al. (2022); Jang et al. (2023).
Despite the crucial role of IFT in the construction of LLMs, there is a significant lack of in-depth research on the mechanisms by which IFT operates. In traditional machine learning, supervised learning aims to fit models to specific tasks and data distributions Goodfellow et al. (2016); Bishop (2006), whereas the impact of IFT on LLMs is markedly different. As shown in Figure 1, on one hand, one of the most apparent effects of IFT is its ability to align the output of LLMs more closely with the latent behavioral norms contained within the IFT data, thereby enabling more effective parameter knowledge expression Zhou et al. (2024); Chen et al. (2023a). On the other hand, many existing studies Li et al. (2023); Cui et al. (2023); Chen et al. (2023b) aim to facilitate domain-specific adaptation of LLMs through IFT, by injecting the world knowledge contained in IFT data into LLMs.
Figure 1: Two potential mechanisms for instruction fine-tuning. 1) Learning, which injects world knowledge in IFT data into LLMs; 2) Self-aligning, which aligns queries with knowledge already in LLMs with similar behavioral norms. Elements with the same color are related.
@@
To this end, this paper designs a knowledge intervention framework for analyzing the underlying mechanisms of IFT. The main idea of our framework is to control the consistency between the knowledge in IFT data and the existing parameter knowledge of LLM, in order to decouple the injection of domain knowledge from the transfer of behavioral norms during IFT. This allows for a separate analysis of the roles of these two crucial factors. Specifically, we first employ in-context learning (ICL) Dong et al. (2022); Brown et al. (2020) to probe the internal parameter knowledge of LLMs. Building on this, we intervene in the composition of existing parameter knowledge and the newly introduced world knowledge within IFT data and then observe the differences in model behavior after IFT using different intervention groups. Based on the framework, we conduct an in-depth analysis to answer the following two critical research questions (RQ):
For RQ1, we initially discover that significant discrepancies between the world knowledge contained in IFT data and the existing parameter knowledge within LLM can substantially undermine the model’s abilities. Performance derived from a set of IFT data that contains incorrect world knowledge but aligns with the model’s parameter knowledge is significantly better than that from a set containing correct world knowledge but inconsistent with the model’s internal parameter knowledge. To dive into this phenomenon, we explicitly supply LLMs with the world knowledge necessary for answering the instruction, integrated into the context. This strategy allows the model to focus on transforming the output behavioral norms instead of jointly learning the inconsistent world knowledge. We discover that the detrimental effects caused by the inconsistency between parameter knowledge and world knowledge can be significantly mitigated by explicitly providing such self-contained IFT data points. These two findings indicate that attempting to introduce world knowledge through IFT that is inconsistent with the model’s parameter knowledge can severely undermine the model, which suggests that the injection of world knowledge does not lie at the center of a successful IFT.
@@
Our experiments reveal the fundamental role of IFT in the construction of LLMs. Essentially, IFT is not a supervised, domain-specific learning process, but a process of self-aligning instruction with existing internal knowledge of LLMs that can be obtained through few-shot in-context learning probing. Our findings not only provide robust theoretical support for the very recently emerged research on self-alignment Sun et al. (2024) and super-alignment Burns et al. (2023), etc., but also shed light on the future direction of data construction, model training, and model evaluation for IFT.
@@ 2 Related Work
By observing output token distribution shift of models before and after IFT, Lin et al. (2024) found that most shifts occur with stylistic tokens, strongly supporting the superficial alignment hypothesis (Zhou et al., 2024), false promise (Gudibande et al., 2023) and related works on IFT data construction Chen et al. (2023a); Shen (2024) and proxy-guided decoding (Liu et al., 2024). While providing intuitive insights, they fall short of providing a comprehensive analysis of IFT’s underlying mechanisms.
Meanwhile, recent efforts have focused on achieving automated alignment, such as self-instruction-tuning Sun et al. (2024); Guo et al. (2024), self-rewarding Yuan et al. (2024) and super-alignment Burns et al. (2023). Despite repeated validations of their effectiveness, there remains limited understanding of their success.
@@ 3 Knowledge Intervention Framework
Specifically, we select four multi-choice datasets from different domains. For each question in each dataset, we employ in-context learning to probe the internal parameter knowledge for each base LLM. Then, we construct multiple instruction datasets by adjusting the consistency of the knowledge within IFT data and model parameter knowledge for each base model. Finally, we analyze the underlying impact of different degrees of world knowledge injection on IFT by fine-tuning LLMs under different settings and examining their performance on the homogeneous, in-domain, and out-of-domain test sets. In the following, we will introduce our knowledge intervention framework in detail.
@@ 3.1 Domain IFT Corpus Setup
In order to facilitate more efficient knowledge consistency identification and IFT evaluation, we select all IFT corpus in the form of multiple choice from four domains: medicine, history, engineering, and jurisprudence. For medicine, we craft a dataset with 20,000 training, 2,206 testing, and 10 development instances by filtering MedMCQA Pal et al. (2022) for entries with explanations and one correct answer, and then applying random sampling. For the other three domains, we procure the relevant items from Xiezhi Benchmark Gu et al. (2024) and held 10 for development and 250 for test  [^2] The availability of multiple-choice datasets tailored to specific domains is extremely limited. Consequently, from this benchmark, we select the three domains with the largest volume of data..
To make a comprehensive evaluation, for each domain’s IFT, we construct three types of test sets: 1) homogeneous test set (HOMO), which is held out from the same multiple-choice dataset of the domain; 2) in-domain test set (ID), including data from MMLU Hendrycks et al. (2021) that belong to the domain; 3) an out-of-domain test set (OOD), comprising instances in MMLU that are from distinct domains. By observing the accuracy performance differences across the three types of test sets, we aim to examine the impact of IFT on various scenarios. Please refer to the Appendix A for more details about our data processing.
@@ 3.2 Parameter Knowledge Probing via Few-shot In-context Learning
Our knowledge intervention framework relies on effectively detecting the parameter knowledge of pre-trained LLMs. To this end, this paper leverages few-shot in-context learning Dong et al. (2022); Brown et al. (2020), which is a widely-used approach for probing the abilities and internal knowledge of pre-trained LLMs Zhang et al. (2023); Wan et al. (2024), to identify the parameter knowledge of our base LLMs. Specifically, we utilize in-context learning to probe the base model’s response to each data item in domain multi-choice dataset and regard the response as the model’s parameter knowledge for this question.
@@ 3.3 Construction of Instruction Data
Upon probing the internal knowledge of pre-trained LLMs, we build instances based on the consistency between world knowledge contained within the domain data and parameter knowledge, thereby constructing different IFT datasets. Specifically, for each domain and base model, we construct IFT data under three settings, including:
  • Harmonious setting, which consists of data where the embedded world knowledge is consistent with model parameter knowledge. This means the pre-trained LLM can answer correctly under in-context learning. In the learning process under this setting, the model only needs to transfer behavioral norms, without the need to acquire additional world knowledge due to the above consistency.
@@
  • Self-aligning setting, which consists of data in which the queries are exactly the same as those in the incompatible set, but we modify the answers corresponding to each query to match the pre-trained LLM’s internal knowledge. Therefore, under this setting, all responses are incorrect, and the model will not learn any additional world knowledge.
To ensure a fair comparison, we maintain the same size across the three groups of data. Meanwhile, to prevent the potential collapse during model training due to the exclusive use of multiple-choice questions, we generate an explanation for the answer and incorporate an equal proportion of general instruction data sampled from alpaca-gpt4-en Peng et al. (2023), thereby ensuring a more stable and real IFT.
@@ 3.4 Experiment Setup
We use LLaMA-2-7B, LLaMA-2-13B, LLaMA-2-70B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023) as the base models of our experiments.
@@
We only calculate loss on outputs, setting epoch to 3, learning rate to $2e^{-5}$ and batch size to 256. For Mistral-7B, we set learning rate to $1e^{-5}$  [^3] Training loss of Mistral-7B using learning rate $2e^{-5}$ does not converge and even spikes.. We use DeepSpeed ZeRO3 Rasley et al. (2020) for LLaMA-2-70B and FSDP Zhao et al. (2023) for the other three. All experiments are implemented on Nvidia A100-80GB GPUs.
@@ 4 Exp-I: Does Learning Domain-specific World Knowledge Matter for IFT?
Table 1: The performance of the four base LLMs after fine-tuning under harmonious (HAR), incompatible (INC), and self-aligning (SELF) settings. For each domain and base model, models fine-tuned on the harmonious dataset and on the self-aligning dataset achieve superior performance compared to those fine-tuned on the incompatible dataset, across all scenarios including homogeneous (HOMO), in-domain (ID), and out-of-domain (OOD) evaluations. The best/second-best performance for each domain and base model in each evaluation is highlighted in bold/underline. The arrows indicate the differences compared to the incompatible setting.
@@
Table 1 reveals that models fine-tuned under the harmonious setting outperform those fine-tuned under the incompatible setting across homogeneous, in-domain, and out-of-domain evaluations [^4] Regarding the evaluations of HOMO, ID, and OOD: The homogeneous, in-domain, and out-of-domain evaluations for each domain cover different subsets and the level of difficulty varies across these subsets. Therefore, the absolute performance differences between HOMO, ID, and OOD are not directly comparable.. Specifically, the harmonious setting yields mean performance gains of 11.27%, 14.58%, and 14.57% over the incompatible setting for homogeneous, in-domain, and out-of-domain tests, respectively. The results indicate that utilizing IFT data, which is consistent with model parameter knowledge and does not inject any additional domain-specific world knowledge, yields superior fine-tuned models. This conclusion holds true across homogeneous, in-domain, and out-of-domain evaluations.
@@
To further investigate the impact of learning domain-specific world knowledge on IFT, we conduct a more direct comparative experiment between the self-aligning and incompatible settings. For each domain and pre-trained LLM, the two settings’ datasets use identical queries that the model cannot answer correctly under in-context learning and have different responses: the incompatible dataset’s responses are correct, reflecting world knowledge, while the self-aligning dataset’s responses represent model parameter knowledge, are incorrect.
@@
The above results demonstrate a significant decline of performance in models fine-tuned using data that contain correct world knowledge but conflict with model parameter knowledge, compared to using consistent IFT data aligned with parameter knowledge. This suggests that introducing additional world knowledge through IFT, in cases where it is inconsistent with the parameter knowledge, may not yield the benefits we anticipate. Therefore, the core role of IFT may lie in facilitating the transfer of behavioral norms, rather than injecting additional domain-specific world knowledge. To further validate this conclusion, in the next section, we design a novel method to decouple the conflict knowledge contained in IFT data and present further analysis of this issue.
@@ 5 Exp-II: IFT with Contextualized Knowledge
To this end, we start with the data of the incompatible setting to construct a dataset with contextualized knowledge. Specifically, given an instruction-answer pair in the incompatible group, we employ GPT-3.5555https://openai.com/ to generate the world knowledge that is required to answer the instruction. The knowledge is then concatenated with the original instruction, as well as the answer, to construct an augmented pair. Finally, we use the constructed data to fine-tune LLMs and compare them with the models fine-tuned with vanilla IFT.
@@
Table 2 shows the results of contextualized knowledge decoupling in three base models  [^6] In this experimental setting, ”vanilla” is equivalent to ”incompatible” group.. From the table, it can be observed that fine-tuning the model with data using explicit contextualized knowledge significantly mitigates the adverse effects caused by inconsistencies between parameter knowledge and world knowledge in IFT data. Compared to vanilla IFT using incompatible data, our method achieves an average improvement of 8.16% on LLaMA-2-7B, 9.48% on LLaMA-2-13B, and 3.98% on Mistral-7B. Except in rare cases, IFT using data with contextualized knowledge can significantly improve the effect of IFT in the incompatible setting across homogeneous, in-domain, and out-of-domain evaluations.
The results indicate that we should not force the model to learn additional inconsistent knowledge when the world knowledge in IFT data is inconsistent with model parameter knowledge. Instead, by decoupling knowledge learning and the transfer of behavioral norms, the problems caused by the above knowledge conflicts can be effectively alleviated. This further verifies our observation in Section 4 that additional world knowledge injection may be ineffective or even harmful for IFT. During IFT, the model should focus on transferring the behavioral norms relying on the existing parameter knowledge and the regularity in the IFT data, rather than learning additional world knowledge. Therefore, for RQ1, we conclude: